### v3

@@ 11 Summary and Conclusions
Our efforts to provide a foundation from which to forecast extreme heat waves were guided by several key decisions. To begin, within the supervised learning traditions used for forecasting, the main response variable used in training was a binary label indicating whether an extreme heat wave had occurred or not. With so many extant definitions of heat waves, we settled on a definition emphasizing a heat dome and responses to extreme heat waves that are often dichotomous: preparations to open public buildings with air conditioning as “cooling centers,” having first responders ready to check on elderly people often living alone, making preparations to close public schools in the afternoons giving ample time for parents to arrange for child care, requiring for water breaks at outdoor work sites, re-allocating staff in emergency rooms to handle increased patient volume, and stocking emergency rooms and urgent care centers with adequate medical supplies (e.g., IV fluids, electrolyte solutions, ice packs, heat-related medication), and much more.
@@
Although our work has been informed by the climate science literature, we make no explicit use of models. Our analyses rest on quasi-experimental designs, a supervised learning algorithm well suited for forecasting, and a genetic algorithm for simple feature engineering. Accurate forecasting is our primary goal. Any explanatory results are a secondary benefit.
The classification in the Pacific Northwest and in the greater Phoenix area proved to be extremely accurate. Although the variable importance plots and partial dependence plots were not identical in the two locations, a case can be made that they are qualitatively similar despite important differences in local climate, topography, research designs and data quality. Still, many more study sites surely are needed, ideally with many more grid cells per site.
By taking genuine forecasting seriously, forecasting uncertainty had to be taken seriously as well. We capitalized on recent developments in nested conformal prediction sets that can provide guarantees of valid results even in finite samples as long as the calibration data are exchangeable. However, the manner in which the data were curated violated exchangeability. Future work can put true forecasting in play by not conditioning on the particular days in which extreme heat waves occured. An illustration, whose details would need to be worked out, could use the AIRS data and a trained classifier perhaps of the sort trained on the PNW data. One could consider each day in turn for the PNW summer months of June, July and August of 2025 (assuming AIRS data availability). For each day, the observed predictor values would be input to the trained classifier. On days when the predicted probability for an extreme heat wave 2 weeks later was sufficiently high, an extreme heat wave would be forecasted. The definition of “high” would depend on the predetermined tradeoff of false positives to false negatives. Conformal prediction sets might be valid as well. There would be much to think through, but no conditioning on a known heat wave day.
@@
Even if our work proves useful, it is at best a start. Among the more important issues going forward is our choice of a two week lag between forecasting precursors and heat wave outcomes. We relied on recent research, but a two week lag is hardly definitive. Our extreme heat wave definition also needs to be revisited. Empirically, it seems to work well, but others might work well too.There are many options already in print. Another significant matter is formulating appropriate populations that are instructive and yet provide better balanced data. We suspect that even if promising populations are proposed, endogenous sampling complications will be ongoing. Extreme heat waves may be increasing, but they are still relatively rare, and “compared to what” questions will remain. Also remaining are research designs and analysis that undertake valid forecasting. Finally, there are many steps between hand-tailored research procedures and production quality forecasting tools that can be routinely used to deliver accurate heat dome warnings sufficiently in advance. We did not address such matters here. That would be at least premature.

### v4

@@ 11 Summary and Conclusions
Our efforts to provide a foundation from which to forecast extreme heat waves were guided by several key decisions. To begin, within the supervised learning traditions used for forecasting, the main response variable used in training was a binary label indicating whether an extreme heat wave had occurred or not. With so many extant definitions of heat waves, we settled on an operationalization emphasizing a heat dome and responses to extreme heat waves that are often dichotomous: preparations to open public buildings with air conditioning as “cooling centers,” having first responders ready to check on elderly people often living alone, making preparations to close public schools in the afternoons giving ample time for parents to arrange for child care, requiring for water breaks at outdoor work sites, re-allocating staff in emergency rooms to handle increased patient volume, and stocking emergency rooms and urgent care centers with adequate medical supplies (e.g., IV fluids, electrolyte solutions, ice packs, heat-related medication), and much more.
@@
Although our work has been informed by the climate science literature, we make no explicit use of models. Our analyses rest on quasi-experimental designs, a supervised learning algorithm well suited for forecasting, and a genetic algorithm for simple feature engineering. Accurate forecasting is our primary long term goal. Any explanatory results are a secondary benefit.
Consistent with our extensive exploration of the data, the classification results in the Pacific Northwest and in the greater Phoenix area proved to be extremely accurate. Although the variable importance plots and partial dependence plots were not identical in the two locations, a case can be made that they are qualitatively similar despite important differences in local climate, topography, research designs and data quality. Still, many more study sites surely are needed, ideally with many more grid cells per site.
By taking genuine forecasting seriously, forecasting uncertainty had to be taken seriously as well. We capitalized thinking for nested conformal prediction sets that can provide guarantees of valid coverage even in finite samples as long as the calibration data are exchangeable. However, the manner in which the data were curated violated exchangeability. Future work can put true forecasting in play by not conditioning on the particular days in which extreme heat waves occurred.
@@
Even if our work proves helpful, it is at best a start. Among the more important issues going forward is our choice of a two week lag between forecasting precursors and heat wave outcomes. We relied on recent research, but a two week lag is hardly definitive. Our extreme heat wave operationalization also needs to be revisited. Empirically, it seems to work well, but others might work well too.There are many options already in print. Another significant matter is formulating appropriate populations that are instructive and yet provide better balanced data. We suspect that even if promising populations are proposed, endogenous sampling complications will be ongoing. Extreme heat waves may be increasing, but they are still relatively rare, and “compared to what” questions will remain. Also remaining are research designs and analysis that undertake valid forecasting. Finally, there are many steps between hand-tailored research procedures and production quality forecasting tools that can be routinely used to deliver accurate heat dome warnings sufficiently in advance. We did not address such matters here. That would be at least premature.